How do you solve an equation with variables on each side such as: 7a + 10 = 2a
Answer by oscargut(2103)

7a+10 = 2a
7a-2a+10 =0
5a+10 =0
5a = -10
a = -2
If you need more help with similar questions you can contact me now at:
mthman@gmail.com
Thanks